8.) Sulfur compounds cause "off-odors" in wine, so winemakers want to know the odor threshold, the lowest concentration of a compound that the human nose can detect. The odor threshold for dimethyl sulfide (DMS) in trained wine tasters is about 25 micrograms per liter of wine (μμg/ l). The untrained noses of consumers may be less sensitive, however.

You want to estimate the mean DMS odor threshold among all students, and you would be satisfied to estimate the mean to within ±±0.25 with 95 % confidence.

The standard deviation of the odor threshold for untrained noses is known to be σσ = 9 micrograms per liter of wine.

How large an SRS of untrained students do you need? _____

9.)

Emissions of sulfur dioxide by industry set off chemical changes in the atmosphere that result in "acid rain." The acidity of liquids is measured by pH on a scale of 0 to 14. Distilled water has pH 7.0, and lower pH values indicate acidity. Normal rain is somewhat acidic, so acid rain is sometimes defined as rainfall with a pH below 5.0. Suppose that pH measurements of rainfall on different days in a Canadian forest follow a Normal distribution with standard deviationσ= 0.5. A sample of n days finds that the mean pH isx= 4.8.

Give a 80 % confidence interval for the mean pH μ when n = 5, n = 15, and n = 40.
n= 5____to____

n = 15____ to ___

n = 40___ to____
Answer #1

8)

Standard deviation = 9

Margin of error = e = 0.25

Confidence level = c = 0.95

z critical value for (1+c)/2 = (1+0.95)/2 = 0.975 is

zc = 1.96 (From statistical table of z values)

Sample size (n) :

n = 4978.714

n = 4979 (Round to nearest integer)

You need SRS of 4979 untrained students